# CAMPBELL & REMIRO-AZÓCAR

### v1

@@ 3.5 Other augmented weighting estimators
In Section 3.4, we proposed augmented estimators that combine the predictions of an unweighted outcome model with weights in a weighted average. Nevertheless, there are other ways of constructing augmented weighting estimators. One approach claimed to be doubly robust consists of G-computation based on the predictions of a weighted outcome model.52, 61 Where the target estimand is the ATC, this involves: (1) estimating weights using the methods described in Section 3.2 and Section 3.3; (2) fitting a weighted model for the conditional outcome expectation to the SAT participants; and (3) marginalizing the outcome predictions of the weighted regression over the external control covariate distribution. The resulting estimator for the mean absolute outcome $\mu_{0}^{1}$ would be:
@@
Such estimators are only doubly robust where the outcome model is a generalized linear model (GLM) with a canonical link function $q(\cdot)$.51, 61, 84, 86 The estimator in Equation 31 and the corresponding ATC estimator would not be doubly robust where the GLM link function is non-canonical,51 or where the outcome model is a Cox proportional hazards model or a parametric survival model in the time-to-event setting.87 Nevertheless, results by Gabriel et al (2024) suggest asymptotic equivalence and similar finite-sample performance to the augmented weighting estimators in Section 3.4 for GLMs with canonical link functions fitted via maximum-likelihood,51 provided that the same weights are used and correct model specification. We note that the target of the investigations by Gabriel et al (2024) is the ATE and the modeling approach to weighting.51
@@ 3.6 Variance estimation
To estimate the variance and construct confidence intervals for $\hat{\mu}_{0}^{1}$ and $\widehat{\textrm{ATC}}$, it is possible to use empirical sandwich-type (“robust”) variance estimators to account for the correlation induced by weighting.25, 38 In the specific context of non-randomized comparisons, such as the externally controlled SATs and unanchored ITCs explored in this article, these estimators have exhibited either under-precision or over-precision for the ATT (or the ATC)88, 89 and under-precision for the ATE.36, 65 This is because most implementations ignore the estimation of the propensity score model or the weights, assuming the weights to be fixed quantities.88, 89
Analytic expressions that incorporate weight estimation could be derived,39, 88, 89 but we propose a practical alternative based on the ordinary non-parametric bootstrap,36 explicitly accounting for uncertainty in the weight estimation. This involves resampling with replacement the concatenated IPD consisting of the SAT and external control data. In each bootstrap iteration, the weight estimation and/or outcome modeling procedures are performed, and $\mu_{0}^{1}$, $\mu_{0}^{0}$ and ATC are re-estimated. Standard errors for $\hat{\mu}_{0}^{1}$, $\hat{\mu}_{0}^{0}$ and $\widehat{\textrm{ATC}}$, are given by the standard deviations across the bootstrap resamples. Subsequently, Wald-type confidence intervals can be constructed. Alternatively, one can directly calculate confidence intervals from the percentiles of the bootstrap resamples, e.g., 2.5% and 97.5% for the 95% confidence interval.
@@ 3.7 External controls with unavailable individual participant data
In the context of unanchored ITCs, the external control is often a historical trial for which IPD are unavailable, due to privacy or confidentiality reasons. In this case, only published AD are available for the external control.25, 28, 32 Such data consist of marginal summary moments $\hat{\mathbf{\theta}}$ from reported tables of baseline characteristics, typically only including means and standard deviations (for continuous covariates), and an estimate $\hat{\mu}_{0}^{0}$ of the mean absolute outcome under the control in the external data source.25, 28, 38, 69, 73 An important shortfall of this scenario is the need to assume that $\hat{\mathbf{\theta}}_{j}=\mathbf{\theta}_{j}$ with zero variability for the covariate balance function moments $j=1,\dots,k$, i.e., that the external control covariate distributional data are fixed.40 While this may be reasonable with large sample sizes for the external control, it can otherwise result in overly precise inferences and inflated Type I error rates.40
In this setting, for all methods except (non-augmented) MAIC, one must simulate $M$ individual-level covariate profiles from the assumed covariate distribution of the external control based on published summary statistics.46, 81, 90 The number $M$ of hypothetical subject profiles should be relatively large, e.g., $M=1000$, to minimize sampling variability and random seed sensitivity, and does not necessarily need to match the original sample size $n_{0}$ of the external control.46, 81, 90 Necessary information to infer the joint covariate distribution of the external control, e.g., distributional forms and correlation structures, is rarely published. Hence, this must be borrowed from other data sources or selected based on theoretical properties, following recommendations in the literature.25, 28, 46, 81, 90
The notation and procedures for Section 3.2, Section 3.4 and Section 3.5 change as follows. The observed IPD for the SAT is stacked with the simulated subject-level covariate data for the external control. The concatenated dataset is now $(S_{i},\mathbf{X}_{i},T_{i},Y_{i})$ for $i=1,\dots,n_{1},n_{1}+1,\dots,n_{1}+M$. For the SAT subjects $i=1,\dots,n_{1}$, we have $S_{i}=1$ and $T_{i}=1$, with $\mathbf{X}_{i}$ and $Y_{i}$ corresponding to the actual covariate and outcome values observed in the trial. For the hypothetical external controls $i=n_{1}+1,\dots,n_{1}+M$, we have $S_{i}=0$ and $T_{i}=0$, the values of $\mathbf{X}_{i}$ are simulated, and $Y_{i}$ are unavailable (but not required for the analysis given that the target estimand is the ATC and that estimate $\hat{\mu}_{0}^{0}$ is available from published results).
The general form of the ATC inverse odds weighting estimators described in Section 3.2 is now:
@@
For the G-computation ATC estimator outlined in Section 3.4, we now have:
@@
The general form of our doubly robust augmented weighting ATC estimators, proposed in Section 3.4, is now:
@@
where $K$ is a constant, $\epsilon_{i}^{1}=Y_{i}-\hat{Y}_{i}^{1}$ and $\hat{v}_{i}$ are a residual term and a weight estimate, respectively, for $i=1,\dots,n_{1}$, and the potential outcome predictions $\hat{Y}_{i}^{1}$ under the active intervention are generated for all SAT subjects and hypothetical external controls $i=1,\dots,n_{1},n_{1}+1,\dots,n_{1}+M$.
For the augmented “weighted G-computation” estimator in Section 3.5, the outcome predictions would be averaged over the simulated covariates for the external control. The resulting estimator for the mean absolute outcome $\mu_{0}^{1}$ is $\hat{\mu}_{0}^{1}=\frac{1}{M}\sum_{i=n_{1}+1}^{n_{1}+M}\hat{Y}_{i}^{1}=\frac{1}{M}\sum_{i=n_{1}+1}^{n_{1}+M}q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}}_{v})\right)$, which would be input into Equation 32 for estimation of the ATC.
SE(ATC^)=($\textrm{SE}\left(\widehat{\textrm{ATC}}\right)=\sqrt{\big{(}\textrm{SE}\left(g(\hat{\mu}^{1}_{0})\right)\big{)}^{2}+\big{(}\textrm{SE}\left(g(\hat{\mu}^{0}_{0})\right)\big{)}^{2}},$ (33)
is used to estimate the standard error of the ATC, where $\textrm{SE}\left(g(\hat{\mu}^{0}_{0})\right)$ is derived from published aggregate-level results.46, 77, 90 A limitation of the above formula is that it assumes that the mean absolute outcomes are statistically independent. Moreover, while computing $\textrm{SE}\left(g(\hat{\mu}^{0}_{0})\right)$ is a trivial exercise for continuous and binary outcomes, (e.g., there is a closed-form formula for the standard error of the log-odds using the Delta method), it can be challenging for other outcomes such as those in the time-to-event setting.77, 90 Once $\textrm{SE}\left(\widehat{\textrm{ATC}}\right)$ is computed, Wald-type confidence intervals can be readily constructed.
3.8 Targeting the average treatment effect in the treated
We briefly adapt the methodologies in Section 3.2 to Section 3.5 so that these target the ATT. We assume that there is full IPD availability and that $\hat{\mu}_{1}^{1}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}Y_{i}$ is consistent for $\mu_{1}^{1}$.
For the modeling-based approach to odds weighting in Section 3.2, external control subjects are weighted by their conditional odds of SAT participation to transport the external control outcomes to the SAT (sub) population. SAT subjects are unweighted and external control subjects $i=n_{1}+1,\dots,n$ are weighted by $\hat{w}_{i}=\hat{e}_{i}/(1-\hat{e}_{i})$. Assuming correct specification of the propensity score model, the estimated weights would balance the covariate distribution of the external control with respect to that of the SAT, enabling consistent estimation of mean absolute outcome $\mu^{0}_{1}$ and the ATT. Propensity score predictions that are close to one lead to extreme weights and imprecise ATT estimation, particularly where the sample size of the external control is small.
A G-computation estimator such like that described in Section 3.4 but for the ATT requires postulating a model for the potential outcome expectation under the control, fitted to the external control participants. Based on the fitted model $m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})$, the potential outcome under the control is predicted for each subject $i=1,\dots,n_{1}$ in the SAT: $\hat{Y}_{i}^{0}=q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})\right)$. The potential outcome predictions are averaged over the empirical covariate distribution of the SAT, resulting in the ATT estimator:
Our doubly robust augmented weighting estimators, proposed in Section 3.4, would target the ATT as follows. Based on an outcome model $m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})$ fitted to the external control participants, the potential outcome under the control is predicted for all subjects $i=1,\dots,n$·in the SAT and the external control: $\hat{Y}_{i}^{0}=q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})\right)$. The potential outcome predictions are augmented with a weighted average of residuals for the external control subjects. The general form of the doubly robust augmented weighting estimators for the ATT is:
$\widehat{\textrm{ATT}}=g\underbrace{\left(\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}Y_{i}\right)}_{\hat{\mu}_{1}^{1}}-g\underbrace{\left(\frac{1}{K}\sum_{i=n_{1}+1}^{n}\hat{v}_{i}\epsilon_{i}^{0}+\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\hat{Y_{i}}^{0}\right)}_{\hat{\mu}_{1}^{0}},$
where $K$ is a constant, $\hat{v}_{i}$ is a weight estimate and $\epsilon_{i}^{0}=Y_{i}-\hat{Y}_{i}^{0}$ is a residual term for subject $i=n_{1}+1,\dots,n$ in the external control.
An augmented “weighted G-computation” estimator akin to that described in Section 3.5 would target the ATT by: (1) estimating weights for the odds of SAT participation; (2) fitting a weighted model $m({\mathbf{X}_{i}};\hat{\mathbf{\beta}}_{v})$ for the conditional outcome expectation to the external control participants; and (3) averaging the outcome predictions of the weighted regression over the SAT covariate distribution. The resulting estimator for the mean absolute outcome $\mu_{1}^{0}$ is $\hat{\mu}_{1}^{0}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\hat{Y}_{i}^{0}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}}_{v})\right)$, which is then substituted into Equation 34 for estimation of the ATT.
@@ 4.1 Aims
We conducted a simulation study to evaluate the performance of various estimators under different conditions. The simulation study design was planned following the structured “ADEMP” approach outlined by Morris et al (2019),91 to ensure reproducibility and meaningful conclusions. Specifically, we clearly defined research aims, data-generating mechanisms under controlled scenarios and estimands, and assessed the performance of several estimators using relevant performance measures: bias, variance and coverage. All simulations and analyses were performed using R statistical software version 4.3.1.92 The files and code required to run the simulations are publicly available on Github at https://github.com/harlanhappydog/DRAWE-.
@@ 4.2 Data-generating mechanisms
We simulated data inspired by the data-generating mechanisms in a simulation study by Kang and Schafer (2007).85 Some modifications were required since Kang and Schafer (2007) considered continuous-valued outcomes,85 while we consider binary outcomes. The simulated data consist of $\{\mathbf{X}_{i},\mathbf{Z}_{i},T_{i},S_{i},Y_{i}\},i=1,\ldots,n\}$, with the control group fully external such that $S_{i}=T_{i}$, with $n_{1}=\sum_{i=1}^{n}{S_{i}}$, and $n_{0}=\sum_{i=1}^{n}{(1-S_{i})}$, as detailed in Section 3.1. While $\mathbf{X}_{i}$ is observed, $\mathbf{Z}_{i}$ is unobserved. To generate the data, $\mathbf{X}_{i}$ is distributed as $\mathrm{Normal}(0,I_{4})$, for $i$ in 1,…,$n$, and $\mathbf{Z}_{i}$ is then obtained by applying the following transformations:
@@
$Z_{i4}$ $=\textrm{scale}((X_{i2}+X_{i4}+20)^{2}).$
where $\textrm{scale}(\dot{)}$ indicates normalization such that $Z_{i1},Z_{i2},Z_{i3}$ and $Z_{i4}$ each have mean of 00 and standard deviation of $1$. Note that these transformations for deriving $\mathbf{Z}_{i}$ from $\mathbf{X}_{i}$ are similar to the ones detailed by Kang and Schafer (2007)85 but not identical, with changes made to highlight the consequences of model misspecification.
@@
    In this scenario, both the outcome model and the propensity score model for data source assignment are correctly specified. The distribution of the covariates is such that overlap across the two groups is relatively high, with overlap proportions of 0.68, 0.85, 0.92, and 0.85 for $X_{1}$,$X_{2}$,$X_{3}$,and $X_{4}$, respectively, see Figure 3 in the Supplementary Material.
@@
    The relevant covariate adjustment approaches would fit an outcome model (of the same form) to the observed $\mathbf{X}_{i}$, as the $\mathbf{Z}_{i}$ used for the true outcome-generating process are unobserved. This implies that the outcome model is incorrectly specified. The propensity score model is still correctly specified. The distribution of the covariates is such that overlap across the two groups is relatively high, with overlap proportions of 0.68, 0.85, 0.92, and 0.85 for $X_{1}$,$X_{2}$,$X_{3}$,and $X_{4}$, respectively, see Figure 3 in the Supplementary Material.
@@
    The relevant covariate adjustment approaches would fit a propensity score model (of the same form) to the observed $\mathbf{X}_{i}$, as the $\mathbf{Z}_{i}$ used for the true data source assignment process are unobserved. This implies that the propensity score model is incorrectly specified. The outcome model is correctly specified. The distribution of the covariates is such that overlap across the two groups is relatively high, with overlap proportions of 0.71, 0.84, 0.99, and 0.89 for $X_{1}$, $X_{2}$, $X_{3}$, and $X_{4}$, respectively, see Figure 4 in the Supplementary Material.
@@
    This scenario implies that both the outcome model and the propensity score model are incorrectly specified. The distribution of the covariates is such that overlap across the two groups is relatively high, with overlap proportions of 0.71, 0.84, 0.99, and 0.89 for $X_{1}$, $X_{2}$, $X_{3}$, and $X_{4}$, respectively, see Figure 4 in the Supplementary Material.

### v2

@@ 3.6 Weighted G-computation
In Section 3.5, we proposed augmented estimators that combine the predictions of an unweighted outcome model with weights in a weighted average. Nevertheless, there are other ways of constructing augmented weighting estimators. One approach popularized by Park et al (2024) park2024unanchored that has been claimed to be doubly robust consists of G-computation based on the predictions of a weighted outcome model.dahabreh2020extending, park2024unanchored Where the target estimand is the ATC, this involves: (1) estimating weights using the methods described in Section 3.2 and Section 3.3; (2) fitting a weighted model for the conditional outcome expectation to the SAT participants; and (3) marginalizing the outcome predictions of the weighted regression over the external control covariate distribution. The resulting estimator for the mean absolute outcome $\mu_{0}^{1}$ would be:
@@
Such estimators are only doubly robust where the outcome model is a generalized linear model (GLM) with a canonical link function $q(\cdot)$.dahabreh2020extending, gabriel2024inverse, robins2007comment, wooldridge2007inverse The estimator in Equation 24 and the corresponding ATC estimator would not be doubly robust where the GLM link function is non-canonical,gabriel2024inverse or where the outcome model is a Cox proportional hazards model or a parametric survival model in the time-to-event setting.gabriel2024propensity Nevertheless, results by Gabriel et al (2024) suggest asymptotic equivalence and similar finite sample performance to the augmented weighting estimators in Section 3.5 for GLMs with canonical link functions fitted via maximum-likelihood,gabriel2024inverse provided that the same weights are used and correct model specification. We note that the target of the investigations by Gabriel et al (2024) is the ATE and the modeling approach to weighting.gabriel2024inverse
@@ 3.7 Variance estimation
To estimate the variance and construct confidence intervals (CIs) for $\hat{\mu}_{0}^{1}$ and $\widehat{\textrm{ATC}}$, it is possible to use empirical sandwich-type (“robust”) variance estimators to account for the correlation induced by weighting.phillippo2016nice, signorovitch2010comparative In the specific context of non-randomized comparisons, such as the externally controlled SATs and unanchored ITCs explored in this article, these estimators have exhibited either under-precision or over-precision for the ATT (or the ATC)reifeis2022variance, kostouraki2024variance and under-precision for the ATE.austin2016variance, hernan2000marginal This is because most implementations ignore the estimation of the propensity score model or the weights, assuming the weights to be fixed quantities.reifeis2022variance, kostouraki2024variance
Analytic expressions that incorporate weight estimation could be derived,cheng2020statistical, reifeis2022variance, kostouraki2024variance but we propose a practical alternative based on the ordinary non-parametric bootstrap,austin2016variance explicitly accounting for uncertainty in the weight estimation. This involves resampling with replacement the concatenated IPD consisting of the SAT and external control data. In each bootstrap iteration, the weight estimation and/or outcome modeling procedures are performed, and $\mu_{0}^{1}$, $\mu_{0}^{0}$ and ATC are re-estimated. Standard errors for $\hat{\mu}_{0}^{1}$, $\hat{\mu}_{0}^{0}$ and $\widehat{\textrm{ATC}}$, are given by the standard deviations across the bootstrap resamples. Subsequently, Wald-type CIs can be constructed. Alternatively, one can directly calculate CIs from the percentiles of the bootstrap resamples, e.g., 2.5% and 97.5% for the 95% CI.
@@ 3.8 External controls with unavailable individual participant data
In the context of unanchored ITCs, the external control is often a historical trial for which IPD are unavailable, due to privacy or confidentiality reasons. In this case, only published AD are available for the external control.phillippo2016nice, phillippo2018methods, lambert2023enriching Such data consist of marginal summary moments $\hat{\mathbf{\theta}}$ from reported tables of baseline characteristics, typically only including means and standard deviations (for continuous covariates), and an estimate $\hat{\mu}_{0}^{0}$ of the mean absolute outcome under the control in the external data source.phillippo2016nice, phillippo2018methods, signorovitch2010comparative, wang2021matching, jackson2021alternative An important shortfall of this scenario is the need to assume that $\hat{\mathbf{\theta}}_{j}=\mathbf{\theta}_{j}$ with zero variability for the covariate balance function moments $j=1,\dots,k$, i.e., that the external control covariate distributional data are fixed.josey2021transporting While this may be reasonable with large sample sizes for the external control, it can otherwise result in overly precise inferences and inflated Type I error rates.josey2021transporting
In this setting, for all methods except (non-augmented) MAIC, one must simulate $M$ individual-level covariate profiles from the assumed covariate distribution of the external control based on published summary statistics.remiro2022parametric, remiro2024transportability, ren2024advancing The number $M$ of hypothetical subject profiles should be relatively large, e.g., $M=1000$, to minimize sampling variability and random seed sensitivity, and does not necessarily need to match the original sample size $n_{0}$ of the external control.remiro2022parametric, remiro2024transportability, ren2024advancing Necessary information to infer the joint covariate distribution of the external control, e.g., distributional forms and correlation structures, is rarely published. Hence, this must be borrowed from other data sources or selected based on theoretical properties, following recommendations in the literature.phillippo2016nice, phillippo2018methods, remiro2022parametric, remiro2024transportability, ren2024advancing
The notation and procedures for Section 3.2, Section 3.4, Section 3.5 and Section 3.6 change as follows. The observed IPD for the SAT is stacked with the simulated subject-level covariate data for the external control. The concatenated dataset is now $(S_{i},\mathbf{X}_{i},T_{i},Y_{i})$ for $i=1,\dots,n_{1},n_{1}+1,\dots,n_{1}+M$. For the SAT subjects $i=1,\dots,n_{1}$, we have $S_{i}=1$ and $T_{i}=1$, with $\mathbf{X}_{i}$ and $Y_{i}$ corresponding to the actual covariate and outcome values observed in the trial. For the hypothetical external controls $i=n_{1}+1,\dots,n_{1}+M$, we have $S_{i}=0$ and $T_{i}=0$, the values of $\mathbf{X}_{i}$ are simulated, and $Y_{i}$ are unavailable but not required for the analysis. This is because the target estimand is the ATC and the outcomes for the external group under the control have already been observed factually, with the mean estimate $\hat{\mu}_{0}^{0}$ available from published results.
The general form of the IOW estimators for the ATC, described in Section 3.2, is now:
@@
For the G-computation estimator for the ATC, outlined in Section 3.4, we now have:
@@
The general form of the DR augmented weighting estimators for the ATC, proposed in Section 3.5, is now:
 $\widehat{\textrm{ ATC}}=g\underbrace{\left(\frac{1}{K}\sum_{i=1}^{n_{1}}\hat{\omega}_{i}\epsilon_{i}^{1}+\frac{1}{M}\sum_{i=n_{1}+1}^{n_{1}+M}\hat{Y_{i}}^{1}\right)}_{\hat{\mu}_{0}^{1}}-g\left(\hat{\mu}_{0}^{0}\right),$
where $K$ is a constant, $\epsilon_{i}^{1}=Y_{i}-\hat{Y}_{i}^{1}$ and $\hat{\omega}_{i}$ are a residual term and a weight estimate, respectively, for $i=1,\dots,n_{1}$, and the potential outcome predictions $\hat{Y}_{i}^{1}$ under the active intervention are generated for all SAT subjects and hypothetical external controls $i=1,\dots,n_{1},n_{1}+1,\dots,n_{1}+M$.
For the weighted G-computation estimator in Section 3.6, the outcome predictions would be averaged over the simulated covariates for the external control. The resulting estimator for the mean absolute outcome $\mu_{0}^{1}$ is $\hat{\mu}_{0}^{1}=\frac{1}{M}\sum_{i=n_{1}+1}^{n_{1}+M}\hat{Y}_{i}^{1}=\frac{1}{M}\sum_{i=n_{1}+1}^{n_{1}+M}q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}}_{v})\right)$, which would be input into Equation 25 for estimation of the ATC.
$\textrm{SE}\left(\widehat{\textrm{ATC}}\right)=\sqrt{\big(\textrm{SE}\left(g(\hat{\mu}^{1}_{0})\right)\big)^{2}+\big(\textrm{SE}\left(g(\hat{\mu}^{0}_{0})\right)\big)^{2}},$ (26)
@@ 4.1 Aims
We conducted a simulation study to evaluate the performance of various estimators under different conditions. The simulation study design was planned following the structured “ADEMP” approach outlined by Morris et al (2019),morris2019using to ensure reproducibility and meaningful conclusions. Specifically, we clearly defined research aims, data-generating mechanisms under controlled scenarios and estimands, and assessed the performance of several estimators using relevant performance measures: bias, empirical standard error and coverage. All simulations and analyses were performed using R statistical software version 4.3.1.r2017r The files and code required to run the simulations are publicly available on Github at https://github.com/harlanhappydog/DRAWE-.
@@ 4.2 Data-generating mechanisms
We simulated data inspired by the data-generating mechanisms in a simulation study by Kang and Schafer (2007).kang2007demystifying Some modifications were required since Kang and Schafer (2007) considered continuous-valued outcomes,kang2007demystifying while we consider binary outcomes. The simulated data consist of $(\mathbf{X}_{i},\mathbf{Z}_{i},T_{i},S_{i},Y_{i})$ for $i=1,\ldots,n$, with the control group fully external such that $S_{i}=T_{i}$, with $n_{1}=\sum_{i=1}^{n}{S_{i}}$, and $n_{0}=\sum_{i=1}^{n}{(1-S_{i})}$, as detailed in Section 3.1. While $\mathbf{X}_{i}$ is observed, $\mathbf{Z}_{i}$ is unobserved. To generate the data, $\mathbf{X}_{i}$ is distributed as $\mathrm{Normal}(0,I_{4})$, for $i$ in 1,…,$n$, and $\mathbf{Z}_{i}$ is then obtained by applying the following transformations:
@@
$Z_{i4}$ $=\textrm{scale}((X_{i2}+X_{i4}+20)^{2}),$
where $\textrm{scale}(\dot{)}$ indicates normalization such that $Z_{i1},Z_{i2},Z_{i3}$ and $Z_{i4}$ each have mean of 0 and standard deviation of $1$, i.e., $Z=\textrm{scale}(f(X))=(f(X)-\textrm{mean}(f(X)))/{\textrm{sd}(f(X))}$. Note that these transformations are similar to the ones detailed by Kang and Schafer (2007)kang2007demystifying but not identical, with changes made to highlight the consequences of model misspecification.
@@
    The distribution of the covariates is such that overlap between the two groups is relatively high, with overlap proportions of 0.68, 0.85, 0.92, and 0.85 for $X_{1}$$X_{2}$$X_{3}$,and $X_{4}$, respectively (see Figure 3 in the Supplementary Material).
@@
    The relevant covariate adjustment approaches would fit an outcome model to the observed $\mathbf{X}_{i}$, as the $\mathbf{Z}_{i}$ used for the true outcome-generating process are unobserved. The distribution of the covariates is such that overlap between the two groups is relatively high, with overlap proportions of 0.68, 0.85, 0.92, and 0.85 for $X_{1}$$X_{2}$$X_{3}$,and $X_{4}$, respectively (see Figure 3 in the Supplementary Material).
@@
    The relevant covariate adjustment approaches would balance or fit a propensity score model to the observed $\mathbf{X}_{i}$, as the $\mathbf{Z}_{i}$ used for the true data source assignment process are unobserved. The distribution of the covariates is such that overlap between the two groups is relatively high, with overlap proportions of 0.71, 0.84, 0.99, and 0.89 for $X_{1}$, $X_{2}$, $X_{3}$, and $X_{4}$, respectively (see Figure 4 in the Supplementary Material).
@@
    The distribution of the covariates is such that overlap between the two groups is relatively high, with overlap proportions of 0.71, 0.84, 0.99, and 0.89 for $X_{1}$, $X_{2}$, $X_{3}$, and $X_{4}$, respectively (see Figure 4 in the Supplementary Material).
We assume that there is unlimited access to subject-level data for the SAT and the external control, such that $(\mathbf{X}_{i},T_{i},S_{i},Y_{i})$ are observed for all $i=1,\ldots,n$.